Radakovich Corporation has provided the following data from its activity-based costing system:

Activity Cost Pools

Total Cost

Total Activity

Assembly

$

1,210,110

57,000

machine-hours

Processing orders

$

68,161

2,370

orders

Inspection

$

153,450

1,980

inspection-hours

The company makes 760 units of product F60N a year, requiring a total of 1,610 machine-hours, 106 orders, and 36 inspection-hours per year. The product's direct materials cost is $59.86 per unit and its direct labor cost is $17.35 per unit. The product sells for $145.40 per unit.

According to the activity-based costing system, the product margin for product F60N is: (Round your intermediate calculations to 2 decimal places.)

$51,824.40

$11,805.54

$14,595.54

$14,854.10

Betterton Corporation uses an activity based costing system to assign overhead costs to products. In the first stage, two overhead costs—equipment depreciation and supervisory expense-are allocated to three activity cost pools—Machining, Order Filling, and Other—based on resource consumption. Data to perform these allocations appear below:

Overhead costs:

Equipment depreciation

$

62,750

Supervisory expense

$

7,000

Distribution of Resource Consumption Across Activity Cost Pools:

Activity Cost Pools

Machining

Order Filling

Other

Equipment depreciation

0.50

0.30

0.20

Supervisory expense

0.10

0.40

0.50


In the second stage, Machining costs are assigned to products using machine-hours (MHs) and Order Filling costs are assigned to products using the number of orders. The costs in the Other activity cost pool are not assigned to products. Activity data for the company's two products follow:

Activity:

MHs (Machining)

Orders (Order Filling)

Product A8

3,900

1,400

Product K2

21,500

4,200

Total

25,400

5,600


How much overhead cost is allocated to the Machining activity cost pool under activity-based costing in the first stage of allocation?

$700

$12,025

$31,375

$32,075

Manton Corporation uses an activity based costing system to assign overhead costs to products. In the first stage, two overhead costs—equipment expense and indirect labor—are allocated to the three activity cost pools—Processing, Supervising, and Other—based on resource consumption. Data to perform these allocations appear below:

Overhead costs:

Equipment expense

$

54,500

Indirect labor

$

9,900

Distribution of Resource Consumption Across Activity Cost Pools:

Activity Cost Pools

Processing

Supervising

Other

Equipment expense

0.60

0.20

0.20

Indirect labor

0.10

0.30

0.60


In the second stage, Processing costs are assigned to products using machine-hours (MHs) and Supervising costs are assigned to products using the number of batches. The costs in the Other activity cost pool are not assigned to products. Activity data for the company's two products follow:

Activity:

MHs (Processing)

Batches (Supervising)

Product Z5

6,670

2,160

Product C8

16,330

1,440

Total

23,000

3,600


The activity rate for the Processing activity cost pool under activity-based costing is closest to:

Noree

$1.46 per MH

$1.42 per MH

$3.42 per MH

$.30 per MH
Answer #1
1) Product margin for F60N
Amount
Selling price (760*145.40) 110504
less:Expenses
direct materials (760*59.86) 45493.6
Direct labor (760*17.35) 13186
overhead
Assembly (1210110/57000)*1610= 34180.3
processing (68161/2370)*106 3048.551
inspection (153450/1980)*36= 2790
total expense 98698.45
product margin 11805.54 answer
2) Machining
Equipment depreciation 62750*50% 31375
Supervisory expense (7000*10%) 700
First stage allocation 32075 answer
3) processing
Equipment expense 54500*60%= 32700
Indirect labor 9900*10% 990
33690
Activity rate = 33690/23000
1.464783
1.46 per machine hours- answer